OpenStudy (anonymous):

Hint: Pay attention to the units of measure. You may have to convert from feet to miles several times in this assignment. You can use 1 mile = 5,280 feet for your conversions. 1. Many people know that the weight of an object varies on different planets, but did you know that the weight of an object on Earth also varies according to the elevation of the object? In particular, the weight of an object follows this equation: , where C is a constant, and r is the distance that the object is from the center of Earth. a. Solve the equation for r. b. Suppose that an object is 100 pounds when it

OpenStudy (anonymous):

r = C/AT mc^2

OpenStudy (anonymous):

could you explain the problem please im not really good at these word problems

OpenStudy (anonymous):

not unless you can provide the whole question. I am not a mind reader. I don't know what "this equation" is unless you tell me.

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

i will do it in parts so you see the whole thing

OpenStudy (anonymous):

Hint: Pay attention to the units of measure. You may have to convert from feet to miles several times in this assignment. You can use 1 mile = 5,280 feet for your conversions. 1. Many people know that the weight of an object varies on different planets, but did you know that the weight of an object on Earth also varies according to the elevation of the object? In particular, the weight of an object follows this equation: , where C is a constant, and r is the distance that the object is from the center of Earth.

OpenStudy (anonymous):

a. Solve the equation for r. b. Suppose that an object is 100 pounds when it is at sea level. Find the value of C that makes the equation true. (Sea level is 3,963 miles from the center of the Earth.)

OpenStudy (anonymous):

c. Use the value of C you found in the previous question to determine how much the object would weigh in i. Death Valley (282 feet below sea level). ii. the top of Mount McKinley (20,320 feet above sea level).

OpenStudy (anonymous):

2. The equation gives the distance, D, in miles that a person can see to the horizon from a height, h, in feet. a. Solve this equation for h. b. Long’s Peak in Rocky Mountain National Park, is 14,255 feet in elevation. How far can you see to the horizon from the top of Long’s Peak? Can you see Cheyenne, Wyoming (about 89 miles away)? Explain your answer.

OpenStudy (anonymous):

i think it has to do with the hint in the beginning

OpenStudy (anonymous):

1. Many people know that the weight of an object varies on different planets, but did you know that the weight of an object on Earth also varies according to the elevation of the object? In particular, the weight of an object follows this equation: , where C is a constant, and r is the distance that the object is from the center of Earth. ^ | | what is this equation ?????!!!!!

OpenStudy (anonymous):

does your problem have an equation in it or not?

OpenStudy (anonymous):

w=Cr^-2

OpenStudy (anonymous):

i did not realize it didnt copy over im sorry

OpenStudy (anonymous):

finally! thank you! now we can proceed.

OpenStudy (anonymous):

solve the equation for r \[w = Cr^{-2} \rightarrow w = \frac {C}{r^{2}} \rightarrow r^2 = \frac C w \rightarrow r = \sqrt{\frac C w}\]

OpenStudy (anonymous):

b. Suppose that an object is 100 pounds when it is at sea level. Find the value of C that makes the equation true. (Sea level is 3,963 miles from the center of the Earth.) w = 100 r = 3963 C = w r^2 C = 100 * 3963^2

OpenStudy (anonymous):

can you take it from there?

OpenStudy (anonymous):

i will work on it and ask if i need more help ty

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

im still not getting this problem, maybe i should explain i have a study sheet with 9 more just like the whole problem i copied out I was hoping to have someone show me step by step on each part so i could work the others out on my own with this as my example its my first day with this stuff and it looks like a foreign language to me right now